OpenStudy (anonymous):

Hi. I need to write an equation of a line that contains the given point and is perpendicular to the given line and I don't know how. The ordered pair is (6,4) and the equation is y=3x-2

OpenStudy (anonymous):

So the first thing you need is a slope that is perpendicular to the original line. What is the slope of the original line?

OpenStudy (anonymous):

I'm not sure. The only numbers that were given to me are the numbers that I gave you.

OpenStudy (anonymous):

hint: When you have a linear equation in the form y = mx + b, the slope (m) is the number in front of the x variable.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Now to find a perpendicular slope, you just take the reciprocal and change the sign: If m is the slope, then \(\large -\frac{1}{m}\) would be the perpendicular slope. So what is the perpendicular slope to 3?

OpenStudy (anonymous):

-(1/3)

OpenStudy (anonymous):

Also correct!

OpenStudy (anonymous):

Now to find a line with a particular slope m that goes through a particular point (h,k) we have the point slope formula: \[y-k = m(x-h)\] In this case we want m to be our perpendicular slope of (-1/3) and our (h,k) point to be (6,4). What will the point slope equation of that look like?

OpenStudy (anonymous):

y-4=-(1/3)(x-6)

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

I'm not sure if that is sufficient for your answer or not, but that is an equation for the line. Sometimes they prefer to have it in y=mx + b form, or some other form.

OpenStudy (anonymous):

how would i put it in y=mx+b form?

OpenStudy (anonymous):

Solve it for y.

OpenStudy (anonymous):

1) distribute the -1/3 to each term in parens 2) add 4 to both sides

OpenStudy (anonymous):

would it be y=-(1/3)x+4?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

You forgot about the (-1/3)(-6)

OpenStudy (anonymous):

\(\large y-4 = -\frac{1}{3}(x-6)\)\[\implies y - 4 = -\frac{1}{3}x- (-\frac{1}{3})(6)\]\[\implies y-4 = -\frac{1}{3}x -(-2)\]\[\implies y-4 = -\frac{1}{3}x + 2\]\[\implies y = -\frac{1}{3}x + 6\]